OpenStudy (anonymous):

sin^(-1)(sin(7pi/6))

OpenStudy (lgbasallote):

hint: (this is a usual scenario but does not happen all the time...but most of the questions in schools follow this principle) \[\huge \sin (\sin^{-1} (x)) = x\] \[\huge \sin^{-1} (\sin (x) ) = x\] does that help?

OpenStudy (lgbasallote):

same thin with \[\huge \cos (\cos^{-1} (x)) = x\] \[\huge \cos^{-1} (\cos (x)) = x\]

OpenStudy (lgbasallote):

or \[\huge \tan (\tan^{-1} (x)) = x\] \[\huge \[\tan^{-1} (\tan (x)) = x\]

OpenStudy (lgbasallote):

i really hope im making sense here =))

OpenStudy (anonymous):

ummm..hold on..

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

so does x just equal 7pi/6?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

GREAT! thanks once again!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

what if its like this: cos(sin^(-1)(x)?

OpenStudy (anonymous):

draw a triangle

OpenStudy (anonymous):

sin^(-1)(x) is the same as sin(theta)= x/1

OpenStudy (anonymous):



OpenStudy (anonymous):

so find the missing side using pythagorean's theorem, then find cos(theta), if im wrong, someone please correct me i havent done a problem like this in a while

OpenStudy (anonymous):

here, i'm gonna start a new post with the exact problem.